# Artificial Inductive Bias for Synthetic Tabular Data Generation in Data-Scarce Scenarios

### v1

@@ Appendix A Appendix
To further investigate the impact of sample size on divergence metric performance, we conducted an additional validation experiment using a reduced number of samples for both $M$ and $L$. Specifically, we set $M=100$ and $L=100$, representing a scenario with limited data availability. As noted in [20], having enough samples is crucial for accurate distribution comparisons. Therefore, we anticipated that this low sample count scenario would yield less reliable divergence results.
The results obtained for this particular validation are presented in Tables 6, 7, 8 and 9. The improvements observed in the previous experiments are not consistent in all datasets under these constrained conditions. Moreover, the divergence values are generally quite small. This suggests that when sample counts are low, the underlying distributions may not be adequately captured, leading to an underestimation of the true disparity between the distributions. Consequently, the divergence may underestimate the actual divergence, resulting in seemingly small divergence values. These findings underscore the importance of having an adequate number of samples when evaluating divergence metrics. With limited data, the reliability of these measures decreases, which can lead to misleading conclusions about similarity between distributions. Therefore, it is essential to consider sample size as a factor when interpreting divergence metric results, particularly in scenarios with limited data availability.
Scenario N M L VAE JS CTGAN JS VAE KL CTGAN KL
Big data 10000 7500 1000 0.079 (0.001) 0.150 (0.002) 0.153 (0.019) 0.420 (0.025)
Low data 300 7500 1000 0.331 (0.004) 0.563 (0.002) 0.697 (0.018) 1.653 (0.015)
Low data 300 100 100 0.030 (0.012) 0.333 (0.027) 0.084 (0.101) 2.197 (0.257)
[…]
Pre-train 300 100 100 -0.002 (0.004) 0.237 (0.014) 0.056 (0.103) 0.978 (0.307)
AVG 300 7500 1000 0.157 (0.004) N/A 0.380 (0.043) N/A
AVG 300 100 100 0.002 (0.002) N/A 0.049 (0.107) N/A
MAML 300 7500 1000 0.300 (0.002) N/A 0.686 (0.037) N/A
MAML 300 100 100 0.002 (0.008) N/A 0.054 (0.116) N/A
DRS 300 7500 1000 0.189 (0.006) N/A 0.427 (0.043) N/A
DRS 300 100 100 0.015 (0.008) N/A 0.089 (0.051) N/A
Table 6: Adult dataset JS and KL results for each Scenario. Big data represents the ideal case where many samples ($N=10,000$) are available to generate reliable synthetic data. Low data represents a more realistic scenario in which a limited number of samples ($N=300$) are available, posing a challenge for synthetic data generation. There are two different Low data rows, depending on the number of samples $M$ and $L$ used for validating. A more reliable case is presented when $M=7,500$ and $L=1,000$ and a less reliable case is presented when $M=100$ and $L=100$. These two cases apply to the following rows too. The next rows compare the divergences obtained by each methodology (pre-training, model averaging, MAML, and DRS) applied to the low data scenario. Results are represented as mean (std). Lower is better. Values in bold indicate an improvement due to the technique used.
Scenario N M L VAE JS CTGAN JS VAE KL CTGAN KL
Big data 10000 7500 1000 0.253 (0.009) 0.463 (0.003) 0.647 (0.045) 1.506 (0.031)
Low data 300 7500 1000 0.840 (0.003) 0.962 (0.002) 4.582 (0.136) 8.994 (0.909)
Low data 300 100 100 0.003 (0.004) 0.482 (0.093) -0.062 (0.113) 2.290 (1.100)
[…]
Pre-train 300 100 100 0.010 (0.006) 0.843 (0.013) 0.061 (0.051) 4.599 (0.629)
AVG 300 7500 1000 0.609 (0.003) N/A 2.596 (0.060) N/A
AVG 300 100 100 -0.001 (0.004) N/A 0.073 (0.088) N/A
MAML 300 7500 1000 0.851 (0.001) N/A 5.176 (0.242) N/A
MAML 300 100 100 0.179 (0.071) N/A 2.256 (0.927) N/A
DRS 300 7500 1000 0.645 (0.006) N/A 2.449 (0.057) N/A
DRS 300 100 100 0.002 (0.004) N/A -0.010 (0.114) N/A
Table 7: News dataset JS and KL results for each Scenario. Big data represents the ideal case where many samples ($N=10,000$) are available to generate reliable synthetic data. Low data represents a more realistic scenario in which a limited number of samples ($N=300$) are available, posing a challenge for synthetic data generation. There are two different Low data rows, depending on the number of samples $M$ and $L$ used for validating. A more reliable case is presented when $M=7,500$ and $L=1,000$ and a less reliable case is presented when $M=100$ and $L=100$. These two cases apply to the following rows too. The next rows compare the divergences obtained by each methodology (pre-training, model averaging, MAML, and DRS) applied to the low data scenario. Results are represented as mean (std). Lower is better. Values in bold indicate an improvement due to the technique used.
Scenario N M L VAE JS CTGAN JS VAE KL CTGAN KL
Big data 10000 7500 1000 0.862 (0.002) 0.777 (0.003) 4.768 (0.072) 3.124 (0.115)
Low data 300 7500 1000 0.927 (0.002) 0.940 (0.003) 13.763 (0.696) 7.470 (0.392)
Low data 300 100 100 0.533 (0.062) 0.682 (0.029) 3.264 (0.555) 3.731 (0.279)
[…]
Pre-train 300 100 100 0.228 (0.018) 0.622 (0.070) 0.698 (0.197) 3.455 (0.426)
AVG 300 7500 1000 0.740 (0.002) N/A 3.489 (0.209) N/A
AVG 300 100 100 -0.001 (0.002) N/A 0.020 (0.108) N/A
MAML 300 7500 1000 0.910 (0.002) N/A 6.436 (0.496) N/A
MAML 300 100 100 0.322 (0.046) N/A 1.030 (0.127) N/A
DRS 300 7500 1000 0.809 (0.003) N/A 4.321 (0.215) N/A
DRS 300 100 100 0.068 (0.006) N/A 0.447 (0.093) N/A
Table 8: King dataset JS and KL results for each Scenario. Big data represents the ideal case where many samples ($N=10,000$) are available to generate reliable synthetic data. Low data represents a more realistic scenario in which a limited number of samples ($N=300$) are available, posing a challenge for synthetic data generation. There are two different Low data rows, depending on the number of samples $M$ and $L$ used for validating. A more reliable case is presented when $M=7,500$ and $L=1,000$ and a less reliable case is presented when $M=100$ and $L=100$. These two cases apply to the following rows too. The next rows compare the divergences obtained by each methodology (pre-training, model averaging, MAML, and DRS) applied to the low data scenario. Results are represented as mean (std). Lower is better. Values in bold indicate an improvement due to the technique used.
Scenario N M L VAE JS CTGAN JS VAE KL CTGAN KL
Big data 10000 7500 1000 0.760 (0.013) 0.531 (0.033) 2.744 (0.084) 2.623 (0.537)
Low data 300 7500 1000 0.920 (0.003) 0.961 (0.002) 6.216 (0.154) 8.841 (0.710)
Low data 300 100 100 0.050 (0.009) 0.681 (0.052) 0.182 (0.150) 4.365 (0.175)
[…]
Pre-train 300 100 100 0.067 (0.004) 0.604 (0.039) 0.286 (0.091) 3.818 (0.189)
AVG 300 7500 1000 0.867 (0.007) N/A 5.798 (0.295) N/A
AVG 300 100 100 0.055 (0.003) N/A 0.167 (0.132) N/A
MAML 300 7500 1000 0.913 (0.003) N/A 6.359 (0.054) N/A
MAML 300 100 100 0.049 (0.002) N/A 0.261 (0.093) N/A
DRS 300 7500 1000 0.835 (0.009) N/A 4.587 (0.166) N/A
DRS 300 100 100 0.066 (0.005) N/A 0.310 (0.173) N/A
Table 9: Intrusion dataset JS and KL results for each Scenario. Big data represents the ideal case where many samples ($N=10,000$) are available to generate reliable synthetic data. Low data represents a more realistic scenario in which a limited number of samples ($N=300$) are available, posing a challenge for synthetic data generation. There are two different Low data rows, depending on the number of samples $M$ and $L$ used for validating. A more reliable case is presented when $M=7,500$ and $L=1,000$ and a less reliable case is presented when $M=100$ and $L=100$. These two cases apply to the following rows too. The next rows compare the divergences obtained by each methodology (pre-training, model averaging, MAML, and DRS) applied to the low data scenario. Results are represented as mean (std). Lower is better. Values in bold indicate an improvement due to the technique used.

### v2

@@ Appendix A Additional Experiments under Constrained Validation Settings
To assess the impact of sample size on divergence metrics, we conducted additional validation using reduced samples for $M$ and $L$, setting $M=100$ and $L=100$ to simulate a low-data scenario. As highlighted in (apellaniz2024synthetic, ), sufficient samples are crucial for accurate distribution comparisons, so we expected less reliable divergence results in this setting.
Scenario N M L DIVERGENCE VALIDATION UTILITY VALIDATION
VAE JS CTGAN JS VAE KL CTGAN KL Real Acc VAE Synth Acc CTGAN Synth Acc
Big data 10000 7500 1000 0.115 (0.002) 0.136 (0.002) 0.272 (0.014) 0.385 (0.025) 0.777 (0.001) 0.748 (0.002) 0.739 (0.006)
Low data 300 100 100 0.113 (0.009) 0.515 (0.011) 0.773 (0.181) 2.694 (0.381) 0.726 (0.006) 0.669 (0.004) 0.675 (0.007)
Pre-train 300 100 100 0.001 (0.001) 0.282 (0.017) 0.037 (0.139) 1.115 (0.265) N/A 0.727 (0.003) 0.692 (0.004)
AVG 300 100 100 0.003 (0.004) N/A 0.085 (0.076) N/A N/A 0.729 (0.004) N/A
MAML 300 100 100 0.105 (0.004) N/A 0.474 (0.099) N/A N/A 0.699 (0.005) N/A
DRS 300 100 100 0.001 (0.001) N/A -0.020 (0.109) N/A N/A 0.726 (0.004) N/A
(a) Adult dataset
Scenario N M L DIVERGENCE VALIDATION UTILITY VALIDATION
VAE JS CTGAN JS VAE KL CTGAN KL Real MSE VAE Synth MSE CTGAN Synth MSE
Big data 10000 7500 1000 0.742 (0.005) 0.746 (0.004) 3.243 (0.152) 2.920 (0.099) 0.325 (0.002) 0.375 (0.000) 0.369 (0.000)
Low data 300 100 100 0.158 (0.060) 0.580 (0.042) 1.171 (0.270) 4.526 (0.699) 0.371 (0.022) 0.520 (0.000) 0.691 (0.000)
Pre-train 300 100 100 0.269 (0.037) 0.694 (0.008) 0.908 (0.065) 2.958 (0.230) N/A 0.446 (0.000) 0.536 (0.000)
AVG 300 100 100 -0.001 (0.008) N/A 0.006 (0.080) N/A N/A 0.363 (0.000) N/A
MAML 300 100 100 0.495 (0.051) N/A 2.516 (0.390) N/A N/A 0.765 (0.000) N/A
DRS 300 100 100 0.250 (0.032) N/A 0.944 (0.209) N/A N/A 0.442 (0.000) N/A
(b) King dataset
Scenario N M L DIVERGENCE VALIDATION UTILITY VALIDATION
VAE JS CTGAN JS VAE KL CTGAN KL Real Acc VAE Synth Acc CTGAN Synth Acc
Big data 10000 7500 1000 0.395 (0.004) 0.635 (0.005) 0.911 (0.027) 1.678 (0.059) 0.861 (0.003) 0.672 (0.007) 0.066 (0.013)
Low data 300 100 100 0.055 (0.035) 0.826 (0.004) 0.978 (0.241) 7.528 (1.065) 0.609 (0.027) 0.401 (0.007) 0.038 (0.004)
Pre-train 300 100 100 0.049 (0.005) 0.619 (0.022) 0.297 (0.101) 4.856 (0.431) N/A 0.392 (0.005) 0.040 (0.013)
AVG 300 100 100 -0.020 (0.017) N/A 0.322 (0.100) N/A N/A 0.405 (0.006) N/A
MAML 300 100 100 0.014 (0.045) N/A 0.845 (0.286) N/A N/A 0.302 (0.008) N/A
DRS 300 100 100 0.007 (0.032) N/A 0.360 (0.090) N/A N/A 0.429 (0.006) N/A
(c) Letter dataset
Table 3: Additional results for each scenario. ‘Big data’ represents the ideal case with ample samples ($N=10,000$) for reliable synthetic data generation, while ‘Low data’ simulates a more constrained setting ($N=300$). The ‘Low data’ rows reflect a less reliable validation case ($M=100$, $L=100$). The next rows compare metrics obtained with different methodologies (‘Pre-train’, ‘AVG’, ‘MAML’, and ‘DRS’) in the ‘Low data’ setting. Results are shown as mean (std), where lower values indicate better performance. For divergence metrics (JS, KL), lower values indicate better performance. For utility metrics, the interpretation depends on the task: for classification tasks (accuracy), higher values are better; for regression tasks (MSE), lower values are better. Bold values denote improvements due to the technique. A slightly negative divergence value close to zero is acceptable, as it reflects an estimation error from the divergence approximation.
Table˜3 shows that, under constrained validation conditions ($M=100$, $L=100$), the improvements observed with our proposed techniques are less consistent across datasets. Divergence values, particularly for JS, remain artificially low, suggesting that small validation sets lack the statistical power to capture meaningful distributional differences. This can lead to misleading conclusions about the similarity between real and synthetic distributions. In contrast, utility validation remains relatively stable across most cases, showing similar trends to those observed with larger $M$ and $L$ values. This indicates that the generative models are able to capture the key informative features needed for classification or regression, even when the global distributional alignment cannot be reliably assessed due to sample scarcity.
Appendix B Additional Datasets
In addition to the three main datasets described in 3.1, we also evaluated our methodology on two supplementary public datasets. These additional datasets provide further evidence of the generalizability of our approach across diverse domains and data modalities. Their inclusion reinforces the robustness of our findings beyond the core evaluation set. Below, we describe each of them:
  • News: The News Popularity Prediction dataset (news, ) includes 39,644 samples and 58 features from articles published on the Mashable news blog. It is a multivariate dataset with continuous and categorical variables that predict article popularity based on social media shares.
  • Intrusion: The KDD Cup 1999 Data (intrusion, ) consists of 494,021 samples and 39 features for classifying connections in a military network environment. It contains categorical, integer, and binary attributes and was used in The Third Knowledge Discovery and Data Mining Competition.
These datasets were selected to diversify the experimental setting in terms of sample size, number of features, and classification difficulty. As with the primary datasets, they were used to test both divergence-based and utility-based evaluations under different data scarcity scenarios. Results for these datasets can be found in Table˜4.
Scenario N M L DIVERGENCE VALIDATION UTILITY VALIDATION
VAE JS CTGAN JS VAE KL CTGAN KL Real Acc VAE Synth Acc CTGAN Synth Acc
Big data 10000 7500 1000 0.223 (0.002) 0.484 (0.010) 0.525 (0.025) 1.432 (0.066) 0.699 (0.007) 0.622 (0.002) 0.616 (0.008)
Low data 300 7500 1000 0.831 (0.004) 0.966 (0.002) 4.301 (0.114) 10.961 (0.317) 0.696 (0.009) 0.564 (0.016) 0.470 (0.030)
Low data 300 100 100 0.182 (0.022) 0.864 (0.017) 0.460 (0.120) 5.673 (0.546) 0.555 (0.043) 0.576 (0.009) 0.512 (0.028)
Pre-train 300 7500 1000 0.734 (0.005) 0.935 (0.001) 3.403 (0.098) 10.441 (0.544) N/A 0.596 (0.003) 0.525 (0.015)
Pre-train 300 100 100 0.009 (0.009) 0.701 (0.010) 0.060 (0.088) 3.868 (0.570) N/A 0.595 (0.004) 0.513 (0.023)
AVG 300 7500 1000 0.618 (0.003) N/A 2.542 (0.054) N/A N/A 0.588 (0.002) N/A
AVG 300 100 100 -0.001 (0.004) N/A -0.101 (0.114) N/A N/A 0.587 (0.002) N/A
MAML 300 7500 1000 0.831 (0.008) N/A 4.381 (0.129) N/A N/A 0.577 (0.008) N/A
MAML 300 100 100 0.002 (0.003) N/A 0.101 (0.132) N/A N/A 0.580 (0.006) N/A
DRS 300 7500 1000 0.603 (0.015) N/A 2.467 (0.050) N/A N/A 0.605 (0.011) N/A
DRS 300 100 100 0.026 (0.006) N/A 0.034 (0.120) N/A N/A 0.609 (0.010) N/A
(a) News dataset
Scenario N M L DIVERGENCE VALIDATION UTILITY VALIDATION
VAE JS CTGAN JS VAE KL CTGAN KL Real Acc VAE Synth Acc CTGAN Synth Acc
Big data 10000 7500 1000 0.795 (0.014) 0.459 (0.002) 3.094 (0.113) 1.724 (0.106) 0.742 (0.369) 0.330 (0.178) 0.590 (0.096)
Low data 300 7500 1000 0.936 (0.003) 0.969 (0.004) 8.297 (0.516) 9.877 (0.770) 0.844 (0.173) 0.567 (0.107) 0.733 (0.008)
Low data 300 100 100 0.228 (0.033) 0.506 (0.021) 1.600 (1.133) 4.353 (0.610) 0.917 (0.010) 0.575 (0.036) 0.735 (0.025)
Pre-train 300 7500 1000 0.797 (0.008) 0.965 (0.001) 4.103 (0.243) 8.409 (0.842) N/A 0.756 (0.022) 0.755 (0.016)
Pre-train 300 100 100 0.066 (0.014) 0.502 (0.044) 0.236 (0.076) 3.680 (0.333) N/A 0.780 (0.011) 0.759 (0.013)
AVG 300 7500 1000 0.850 (0.002) N/A 4.934 (0.134) N/A N/A 0.801 (0.007) N/A
AVG 300 100 100 0.058 (0.003) N/A 0.330 (0.114) N/A N/A 0.798 (0.006) N/A
MAML 300 7500 1000 0.945 (0.005) N/A 8.046 (0.457) N/A N/A 0.652 (0.061) N/A
MAML 300 100 100 0.368 (0.023) N/A 3.131 (0.555) N/A N/A 0.532 (0.145) N/A
DRS 300 7500 1000 0.844 (0.015) N/A 5.613 (0.435) N/A N/A 0.544 (0.242) N/A
DRS 300 100 100 0.043 (0.011) N/A 0.289 (0.187) N/A N/A 0.585 (0.121) N/A
(b) Intrusion dataset
Table 4: Results for additional datasets in each scenario. ‘Big data’ represents the ideal case with ample samples ($N=10,000$) for reliable synthetic data generation, while ‘Low data’ simulates a more constrained setting ($N=300$). Two ‘Low data’ rows reflect different validation sample sizes: a more reliable case ($M=7,500$, $L=1,000$) and a less reliable case ($M=100$, $L=100$), which apply to the following rows. The next rows compare metrics obtained with different methodologies (‘Pre-train,’ ‘AVG,’ ‘MAML,’ and ‘DRS’) in the ‘Low data’ setting. Results are shown as mean (std), where lower values indicate better performance. For divergence metrics (JS, KL), lower values indicate better performance. For utility metrics, the interpretation depends on the task: for classification tasks (accuracy), higher values are better; for regression tasks (MSE), lower values are better. Bold values denote improvements due to the technique. A slightly negative divergence value close to zero is acceptable, as it reflects an estimation error from the divergence approximation.
Table˜4 confirms the generalizability of our findings across the additional News and Intrusion datasets. Consistent with the main experimental results, the VAE model outperforms CTGAN in both divergence and utility metrics across most scenarios. Furthermore, the proposed techniques—particularly ‘Pre-train,’ ‘AVG,’ and ‘DRS’—consistently improve over the baseline (‘Low data’) in terms of JS and KL divergence, as well as downstream performance (accuracy). The only exception remains ‘MAML,’ which shows less stable behavior and, in several cases, fails to provide improvements. These results reinforce the robustness of our methodology and its applicability to diverse tabular domains.
Appendix C Computational Resources and Training Time Analysis
All experiments were conducted on a prosumer-grade workstation equipped with an AMD Ryzen Threadripper PRO 5975WX processor featuring 32 cores and 64 threads. This CPU operates at a base frequency of 3.6 GHz with a boost up to 4.5 GHz, offering substantial parallel computing capacity. The machine was configured with 128 GB of DDR4 RAM and a 1 TB NVMe SSD for fast data access, along with additional storage on dual 4 TB HDDs. Although the workstation included an NVIDIA RTX 4090 graphics card, GPU acceleration was not required for training the VAE-based model, as all experiments involving the VAE were performed on the CPU. On the other hand, the CTGAN implementation used in this work utilizes GPU acceleration by default.
Despite the added complexity introduced by some of the proposed techniques, training times remained practical for both models. Table˜5 summarizes the average training durations for each dataset and methodology over 10 runs using different random seeds. The VAE, although trained with 10 different seeds, achieved competitive training times. Even in the most demanding scenario (‘Big data’), the average training time across datasets remained below 37 minutes (2,198 seconds), and no single experiment exceeded one hour. CTGAN experiments were significantly faster, especially due to GPU usage, yet both architectures can be considered efficient and lightweight in practice. Importantly, for both models, applying the proposed methodologies in the low-data regime, such as ‘Pre-train,’ ‘AVG,’ ‘MAML,’ and ‘DRS,’ results in only moderate increases in training time compared to the baseline ‘Low data’ scenario. This demonstrates the practicality and scalability of our approach: the benefits of inductive bias integration are achieved with minimal additional computational cost, making the methods suitable for real-world settings where time and resources are constrained.
Dataset Big data Low data Pre-train AVG MAML DRS
VAE CTGAN VAE CTGAN VAE CTGAN VAE CTGAN VAE CTGAN VAE CTGAN
Adult 1893.25 382.356 61.6617 19.0268 1367.84 402.089 145.509 125.314 463.661
News 7248.41 532.651 115.353 27.1911 4593.78 558.03 139.851 534.483 386.022
King 1282.78 393.782 70.4384 20.6333 567.242 416.421 75.209 204.017 133.523
Intrusion 1651.26 570.559 82.2507 29.3055 4909.34 604.459 187.802 234.104 923.386
Letter 821.65 299.236 96.9268 16.5833 776.11 323.648 84.3122 193.772 210.785
Average 2579.87 435.317 85.3269 22.94798 2442.06 460.929 126.537 258.338 423.875
Table 5: Average execution times (in seconds) across 10 runs with different random seeds for the various scenarios and datasets.
Overall, this analysis demonstrates that the benefits provided by inductive bias techniques come at a relatively low computational cost. Even when training multiple VAE instances or applying more complex optimization procedures, the observed increases in runtime remain within practical limits. This reinforces the applicability of our methods in real-world, resource-constrained environments and supports their use in privacy-sensitive domains where synthetic data are needed, but GPU availability may be limited.